At the beginning of the walk, robert and juan are 5.5 miles apart. If they leave at the same time and walk in the same direction, robert overtakes juan in 11 hours. If they walk towrds eachother, they meet in 1 hour. What are their speeds?
Answer by mananth(16949)

At the beginning of the walk, robert and juan are 5.5 miles apart. If they leave at the same time and walk in the same direction, robert overtakes juan in 11 hours. If they walk towrds eachother, they meet in 1 hour. What are their speeds?
..
let Robert's speed be x mph
Juans speed be y
if they walk in same direction their effective speed will be x-y
time they take to meet = 5.5/x-y =11
5.5=11(x-y)
5.5 =11x-11y
11x-11y=5.5............1
when they walk towards each other their speed will be x+y mph
5.5/(x+y)= 1
x+y=5.5..............2
multiply 2 by 11
11x+11y=60.5
add it to equation 1
22x=66
/22
x= 3mph speed of Robert
2.5 mph = speed of Juan
...
m.ananth@hotmail.ca